In September 2016, the management of Rye Company assembles the following
data in preparation of budgeted merchandise purchases for the months of October and November.
1) Expected Sales
a) October $1,500,000
b) November 2,100,000
c) December 2,700,000
2) Cost of goods sold is expected to be 70% of sales
3) Desired ending merchandise inventory is 20% of the next month's cost of goods sold.
4) The beginning inventory at october 1 will be the desired amount.
INSTRUCCIONS
Compute the budgeted merchandise purchases for October and November.
Use a columnar format with separate columns for each month.
Answer #1

Compute purchase

October November
Sales 1500000 2100000
Cost of goods sold 1050000 1470000
Add: Desired ending inventory 1470000*20% = 294000 378000
Total 1344000 1848000
Less: Beginning inventory 1050000*20% = -210000 -294000
Purchase 1134000 1554000